OpenStudy (anonymous):

Find the derivative of f(x) = -2x^2 + 11x at x = 9

OpenStudy (bibby):

can you find the derivative of -2x^2?

OpenStudy (anonymous):

4 @bibby

OpenStudy (bibby):

so the power rule states that \(\large \dfrac{d}{dx}x^n=nx^{n-1}\), right? \(\dfrac{d}{dx}-2x^{2}=-2\dfrac{d}{dx}=x^2\\=-2(2x^{2-1})=-2(2x^1)=-2(2x)=-4x\) do the same thing for the other term

OpenStudy (anonymous):

@bibby that's what I don't get

OpenStudy (bibby):

be more specific, you don't get the power rule?

OpenStudy (anonymous):

how to do it for 11x. and what to do with the 9

OpenStudy (bibby):

when they ask you to find the derivative at x=9, they want you to find \(f'(x)\) and plug in x=9, so \(f'(x)\) is the derivative and \(f'(9)\) is the value at x=9 just rewrite 11x as 11x^1 and then plug it into the power rule

OpenStudy (anonymous):

d/dx + 11x^1 = 11 d/dx = x^1 11(11x ^1-1) = 11(11x) Is this correct so far @bibby

OpenStudy (bibby):

close 1-1=0 so we have \(11x^1\iff 11*1x^{1-1}=11*1*x^0=11*1*1=11\)

OpenStudy (anonymous):

Okay and now I have 4x and 11 so I plug nine in for x then that is my answer? @bibby

OpenStudy (bibby):

f'(x)=-4x+11

OpenStudy (bibby):

yeah

OpenStudy (anonymous):

Thank you so much @bibby